# Can GitHub Issues Help in App Review Classifications?

### v2

@@ Abstract
Keywords App Review Classification  $\cdot$ GitHub Issues  $\cdot$ Intention Mining  $\cdot$ Machine learning.
1 Introduction
@@
App stores contain a large volume of mobile apps related to various fields, each offering different features [1, 2, 3]. Users can express their feedback and experiences within these stores by posting reviews [4, 1]. User reviews contain valuable information, such as bug reports, feature suggestions, requests for improving current features, etc., which can assist developers in future maintenance and development plans [1, 5, 4, 2]. Due to the substantial number of user reviews, particularly in popular apps, various approaches have been presented to automatically classify valuable reviews from the app developer’s point of view [6, 2, 4, 7]. Previous studies often focus on providing novel approaches based on Natural Language Processing (NLP) and Machine Learning (ML) techniques to increase the accuracy of review classification [8, 7] and utilize manually labeled datasets to train and evaluate classification models [7]. They usually prepare these datasets by following a systematic process to minimize human errors in labeling reviews, but this can be both expensive and time-consuming [6, 8]. As reported by Guzman et al. [6], labelers spend an average of 22 hours analyzing 1820 reviews manually. Therefore, it is challenging to prepare a precisely labeled dataset, and thus, they have a relatively low volume and are selected from a small number of apps [8]. Dabrowski et al. [8] identified this problem as a “significant threat” for assessing the generalizability of existing approaches. Due to this limitation, although models trained on a specific dataset perform well during testing on the same dataset, a recent study demonstrates that these models perform poorly when applied to unseen datasets [7]. Devine et al. [7] showed that the performance of transformer-based models, trained separately on seven datasets, decreases when tested on each other. Therefore, the trained models are less generalizable than expected. They raise an important question: “How informative are the predictions of these models when used in the real world?” To address this, they demonstrate that better results can be achieved by combining multiple labeled datasets. Although this approach enhances the performance of the classification models, their results indicate that further improvements in generalizability are still necessary.
@@
However, the challenges of preparing labeled datasets have received comparatively less attention. In some studies, researchers have explored the use of semi-supervised [9] or active learning [10] methods or combined dataset [7] to address this critical challenge [8]. The main problem with these approaches is their reliance on either insufficient existing datasets or human-dependent methods for collecting new data, both of which present significant difficulties as mentioned earlier. So, despite these efforts, the trained models still exhibit low accuracy and require further improvement. To answer this problem, we propose an approach to enhance the generalizability of trained models by augmenting the manually labeled review dataset.
@@
In addition to app stores, some platforms such as Twitter, forums (e.g. VLC and Firefox forums [11]), issue-tracking systems (e.g. GitHub [12]), etc., receive user feedback [7, 13]. Issue-tracking systems are not exclusive to developers, and users can also provide feedback, such as bug reports, feature requests, etc., by creating issues in app repositories [14]. Developers usually assign one or more labels to an issue to present its purpose, topic, or priority [15]. Also, issue bodies have similarities with app store user reviews [16, 17]. Al-Safoury et al. [16] leverage the similarity between issue reports and user reviews to complete mentioned change requests. Also, Zhang et al. [17] utilize this similarity to assist in classifying unlabeled GitHub issues. In comparison, we use issue reports to improve review classifiers by applying linguistic patterns to adapt them to user reviews. Although issues contain more technical details than reviews, they can enrich review-based labeled datasets. In this paper, we augment the manually labeled dataset by integrating processed labeled issues with them to train more generalizable models for review classification. To the best of our knowledge, this paper is the first study that uses labeled issues in GitHub to enhance the generalizability of app review classifiers.
@@
Since the issue tracking system serves as a platform for developing projects, developers are more inclined to contribute to repositories rather than users. Therefore, issue bodies can provide information that may not be visible in app store user reviews, such as technical details, code snippets, error messages, etc. We extract the target information from issue bodies to adapt them to user reviews. We define 19 linguistic patterns, derived from manual analysis of 577 related issue templates and 2,098 main constituent section titles of issue bodies that do not follow defined templates (Section 3.2.2). By processing the issue bodies this way, we bring them closer to the user reviews posted in app stores. Next, we integrate the labeled review datasets with a subset of processed issues selected through three methods Within-App (using processed issues from the same app), Within-Context (using processed issues from similar apps), and Between-App (using randomly selected processed issues) Analysis (Section 3.4). Finally, we define this problem as two binary classification tasks: determining whether a review is a Bug Report and whether a review is a Feature Request. It is possible for a review to fall into both categories, resulting in positive outputs from both classifiers. We fine-tune transformer-based classifiers, which include pre-trained BERT [18], RoBERTa [19], DistilBERT [20], and ALBERT [21] models. These models have demonstrated state-of-the-art performance in previous studies [22, 7] for review classification tasks (Section 3.5).
@@
The results of our experiments demonstrate that augmented datasets can lead to a significant improvement in model performance. The F1-score increased to 6.3 and 7.2 for bug reports and feature requests, respectively. In augmented models related to bug reports, we observe that recall increases up to 9.2 compared to the baseline model, along with an improvement in precision. Notably, recall is particularly crucial for developers in review classifiers, especially when predicting bug reports [1]. A higher recall ensures that less target information is lost. So, our approach enhances the reliability of review classifiers. As expected, the Within-App and Within-Context methods, which are used to train app-specific classifiers, exhibit more improvement compared to the Between-App method, which is used to train the general-purpose classifiers. This outcome can be attributed to the fact that Within-App and Within-Context methods leverage more relevant and app-specific information, leading to better performance.
@@
2 A Motivating Example
(a) An Example of a GitHub Issue [23]
(b) An Example of a Goggle Play Review [24]
Figure 1: Similarity Between Issues and Reviews
Figure 1b shows a Google Play user review [16] for the AtennaPod [24] open-source app, a podcast manager. The main parts of a user review are the body and rating. The review body contains a textual description of the user feedback about a bug encountered, a new feature requested, etc., expressed without a formal structure [6]. For example, in Figure 1b, a user describes the problem of resuming a podcast after the app disconnects from the car.
A user in this review reported a bug informally. However, this bug has been reported more formally in GitHub issues of the AtennaPod repository [23]. Figure 1a shows an issue related to the same bug with the label “type: possible bug”. The main parts of an issue include the title, body, and label. As we can see, a description of the same bug mentioned in the user review (Figure 1b) is provided in the Current Behavior section of the body with slightly different constituent words. The issue title also briefly describes this problem. In this issue, other sections provide details for reproducing the bug, which is unrelated to our purpose. Therefore, although GitHub issues more technically report encountered bugs or describe requested features, they can be used for detecting the review’s intention by extracting target information.
3 Proposed Approach
@@
Figure 2: An Approach Overview
3.1 Data Collection
We first collect data needed for subsequent steps of our approach. Therefore, we investigate the Android repositories available on GitHub to identify appropriate ones that (i) contain adequate labelled issues and (ii) provide relevant and accurate data for creating the auxiliary dataset. Liu et al. [25] provided the AndroZooOpen dataset, which includes information about more than 46K Android repositories. To our best knowledge, this dataset is the largest available collection containing GitHub repositories related to open-source Android apps. This paper uses the AndroZooOpen dataset to find appropriate repositories by examining extracted information.
In January 2023, approximately 43K repositories of this dataset were available on GitHub. We updated important fields, such as the contributor and star counts, to make our investigation more accurate and get labeled issues using the GitHub REST API [26]. Most available repositories had no labeled issue, and only 5,258 (12.1%) repositories contained at least one. Figure 3 shows the repository counts for each number of labeled issues in these 5,258 repositories. As can be seen, most of the repositories have a few labeled issues. Although 20% of them contain more than 30 labeled issues, they constitute 92.3% of collected issues. Therefore, repositories with less than 30 labeled issues provide little information. In order to increase the quality and performance of the auxiliary dataset, we eliminate them for two reasons.
First, in preparing the auxiliary dataset, we ignore unrelated labels to our review intentions (Section 3.2.1) and then identify and extract the target information from selected issues (Section 3.2.2). As explained, more than two-thirds of the collected issues, which are irrelevant to our intentions or do not provide sufficient information, are eliminated in these steps. The presence of many repositories with few labeled issues that ultimately provide little helpful information can reduce the performance of finding similar apps (Section 3.4.1). Second, Figure 4 shows the distribution of the number of contributors and stars for repositories with at least one labeled issue. HLI (High-Labeled Issue) and LLI (Low-Labeled Issue) denote repositories with a minimum and a maximum of 30 labeled issues, respectively. As can be seen, more developers are involved in HLI repositories, which causes more monitoring. Also, developers have given HLI repositories more stars. Considering stars as a measure of repository popularity, following Borges et al. [27], we can conclude that HLI repositories have garnered more interest from developers. Therefore, data provided by HLI repositories with more monitoring and popularity can be more reliable.
@@ 3.1 Data Collection
Table 1: Preliminary Collected Data
@@
Figure 3: Repository Counts per Each Number of Labelled Issues (Repos=Repositories)
(a) Distribution of Stars
(b) Distribution of Contributors
Figure 4: The Comparison of HLI and LLI Repositories
3.2 Preparing the Auxiliary Dataset
@@
Previous studies have classified user reviews according to their intentions [8]. They have focused on different perspectives and granularity to assist developers in maintenance tasks [4, 28]. Table 2 displays the target intention classes of reviews in several studies with available labeled datasets. As seen, two intentions, namely Bug Report and Feature Request, have been repeatedly used with slight differences in their naming. Therefore, given the available data required to create the primary dataset, we follow Devine et al. [7] and consider the three most frequent classes of Bug Report, Feature Request, and Other as our target review intentions.
Table 2: Target Intention Classes in Previous Studies (*For studies that presented multi-level taxonomy, we only mentioned the first level.)
@@
Pagano et al. [29] community, requirement, rating, user experience *
Maalej et al. [1] bug report, feature request, user experience, rating
Panichella et al. [30] information giving, information seeking, feature request, problem discovery, other
Gu et al. [31] aspect evaluation, bug reports, feature requests, praise and others
Guzman et al. [6] bug report, feature strength, feature shortcoming, user request, praise, complaint, usage scenario
Ciurumelea et al. [32] compatibility, usage, resources, pricing, and protection*
Jha et al. [33] bug report, feature request, other
Scalabrino et al. [34] functional bug report, suggestion for new feature, report of performance problems, report of security issues, report of excessive energy consumption, request for usability improvements and other
3.2.1 Filtering the Issue Labels
In GitHub, developers assign labels to an issue based on various criteria, such as intentions (e.g., bug, enhancement), states (e.g., duplicate, invalid), development stages (e.g., implementation, design), app components or features (e.g., component: client, feature: history), priorities (e.g., low priority, high priority), etc. These labels are helpful in managing and organizing issues more efficiently. By default, GitHub provides seven labels: Bug, Enhancement, Documentation, Duplicate, Good First Issue, Help Wanted, Invalid, Question, and Wontfix. Developers also have the option to create new labels for categorizing their issues, which we refer to as user-defined labels. There are no specific rules for defining these labels [15]. As a result, they can be heterogeneous and diverse. Different labels may represent a similar concept. For instance, labels like Type: Enhancement, Enhancement/Feature, Feature-Request are used in many repositories to represent the same concept as the default label Enhancement. Due to this variability, Izadi et al. [15] manually analyzed the issue labels according to their goals. In the previous phase, we collected 318K labeled issues, comprising 5,641 distinct labels. Figure 5 displays the 20 most frequently occurring labels in our preliminary dataset. Similar labels may vary in form (e.g., bug, Bug) or constituent words (e.g., enhancement, feature request). To address this, we first preprocess the labels to unify their forms and remove noisy characters. Then, we manually identify the semantically relevant labels aligned with our review intentions.
Figure 5: The Most Frequent Labels in Preliminary Dataset
@@ 3.2.1 Filtering the Issue Labels
  4. 4.
    Stemming using the SnowballStemmer function in the NLTK library [35]: For example, three labels reproduced, reproduction, and reproducible, which refer to the same concept, are converted to a single term reproduc.
  5. 5.
@@
Table 3: Samples of the Provided Information for Each Preprocessed (Pre) Label
@@
The app crashes, either the foreground or background. **To Reproduce** Steps to reproduce the behavior: 1. Go to Create a new list 2. Click on Ok in the list name dialog, important step 3. Rotate the app
@@
Table 4: Some Preprocessed Labels Related to Target Review Intentions
@@
3.2.2 Extracting Target Information

### v3

@@ Abstract.
App Review Classification, GitHub Issues, Intention Mining, Machine learning
copyright: acmcopyrightjournalyear: 2023journal: TOSEMccs: Software and its engineering Software libraries and repositoriesccs: Information systems Information extraction
1. Introduction
@@
App stores contain a large volume of mobile apps related to various fields, each offering different features (Maalej et al., 2016; Aslam et al., 2020; Abedini et al., 2024). Users can express their feedback and experiences within these stores by posting reviews (Villarroel et al., 2016; Maalej et al., 2016). User reviews contain valuable information, such as bug reports, feature suggestions, requests for improving current features, etc., which can assist developers in future maintenance and development plans (Maalej et al., 2016; Chen et al., 2014; Villarroel et al., 2016; Aslam et al., 2020). Due to the substantial number of user reviews, particularly in popular apps, various approaches have been presented to automatically classify valuable reviews from the app developer’s point of view (Guzman et al., 2015; Aslam et al., 2020; Villarroel et al., 2016; Devine et al., 2023). Previous studies often focus on providing novel approaches based on Natural Language Processing (NLP) and Machine Learning (ML) techniques to increase the accuracy of review classification (Dąbrowski et al., 2022; Devine et al., 2023) and utilize manually labeled datasets to train and evaluate classification models (Devine et al., 2023). They usually prepare these datasets by following a systematic process to minimize human errors in labeling reviews, but this can be both expensive and time-consuming (Guzman et al., 2015; Dąbrowski et al., 2022). As reported by Guzman et al. (Guzman et al., 2015), labelers spend an average of 22 hours analyzing 1820 reviews manually. Therefore, it is challenging to prepare a precisely labeled dataset, and thus, they have a relatively low volume and are selected from a small number of apps (Dąbrowski et al., 2022). Dabrowski et al. (Dąbrowski et al., 2022) identified this problem as a “significant threat” for assessing the generalizability of existing approaches. Due to this limitation, although models trained on a specific dataset perform well during testing on the same dataset, a recent study demonstrates that these models perform poorly when applied to unseen datasets (Devine et al., 2023). Devine et al. (Devine et al., 2023) showed that the performance of transformer-based models, trained separately on seven datasets, decreases when tested on each other. Therefore, the trained models are less generalizable than expected. They raise an important question: “How informative are the predictions of these models when used in the real world?” To address this, they demonstrate that better results can be achieved by combining multiple labeled datasets. Although this approach enhances the performance of the classification models, their results indicate that further improvements in generalizability are still necessary.
@@
However, the challenges of preparing labeled datasets have received comparatively less attention. In some studies, researchers have explored the use of semi-supervised (Deocadez et al., 2017) or active learning (Dhinakaran et al., 2018) methods or combined dataset (Devine et al., 2023) to address this critical challenge (Dąbrowski et al., 2022). The main problem with these approaches is their reliance on either insufficient existing datasets or human-dependent methods for collecting new data, both of which present significant difficulties as mentioned earlier. So, despite these efforts, the trained models still exhibit low accuracy and require further improvement. To answer this problem, we propose an approach to enhance the generalizability of trained models by augmenting the manually labeled review dataset.
@@
In addition to app stores, some platforms such as Twitter, forums (e.g. VLC and Firefox forums (Tizard et al., 2019)), issue-tracking systems (e.g. GitHub (Khalajzadeh et al., 2022)), etc., receive user feedback (Devine et al., 2023; Nayebi et al., 2018). Issue-tracking systems are not exclusive to developers, and users can also provide feedback, such as bug reports, feature requests, etc., by creating issues in app repositories (Mazrae et al., 2021). Developers usually assign one or more labels to an issue to present its purpose, topic, or priority (Izadi et al., 2022). Also, issue bodies have similarities with app store user reviews (Al-Safoury et al., 2022; Zhang et al., 2019). Al-Safoury et al. (Al-Safoury et al., 2022) leverage the similarity between issue reports and user reviews to complete mentioned change requests. Also, Zhang et al. (Zhang et al., 2019) utilize this similarity to assist in classifying unlabeled GitHub issues. In comparison, we use issue reports to improve review classifiers by applying linguistic patterns to adapt them to user reviews. Although issues contain more technical details than reviews, they can enrich review-based labeled datasets. In this paper, we augment the manually labeled dataset by integrating processed labeled issues with them to train more generalizable models for review classification. To the best of our knowledge, this paper is the first study that uses labeled issues in GitHub to enhance the generalizability of app review classifiers.
@@
Since the issue tracking system serves as a platform for developing projects, developers are more inclined to contribute to repositories rather than users. Therefore, issue bodies can provide information that may not be visible in app store user reviews, such as technical details, code snippets, error messages, etc. We extract the target information from issue bodies to adapt them to user reviews. We define 19 linguistic patterns, derived from manual analysis of 577 related issue templates and 2,098 main constituent section titles of issue bodies that do not follow defined templates (Section 3.2.2). By processing the issue bodies this way, we bring them closer to the user reviews posted in app stores. Next, we integrate the labeled review datasets with a subset of processed issues selected through three methods Within-App (using processed issues from the same app), Within-Context (using processed issues from similar apps), and Between-App (using randomly selected processed issues) Analysis (Section 3.4). Finally, we define this problem as two binary classification tasks: determining whether a review is a Bug Report and whether a review is a Feature Request. It is possible for a review to fall into both categories, resulting in positive outputs from both classifiers. We fine-tune transformer-based classifiers, which include pre-trained BERT (ber, 2023), RoBERTa (rob, 2023), DistilBERT (dis, 2023), and ALBERT (alb, 2023) models. These models have demonstrated state-of-the-art performance in previous studies (Hadi and Fard, 2021; Devine et al., 2023) for review classification tasks (Section 3.5).
@@
The results of our experiments demonstrate that augmented datasets can lead to a significant improvement in model performance. The F1-score increased to 6.3 and 7.2 for bug reports and feature requests, respectively. In augmented models related to bug reports, we observe that recall increases up to 9.2 compared to the baseline model, along with an improvement in precision. Notably, recall is particularly crucial for developers in review classifiers, especially when predicting bug reports (Maalej et al., 2016). A higher recall ensures that less target information is lost. So, our approach enhances the reliability of review classifiers. As expected, the Within-App and Within-Context methods, which are used to train app-specific classifiers, exhibit more improvement compared to the Between-App method, which is used to train the general-purpose classifiers. This outcome can be attributed to the fact that Within-App and Within-Context methods leverage more relevant and app-specific information, leading to better performance.
@@
2. A Motivating Example
Figure 1. Similarity Between Issues and Reviews
Figure 1(b) shows a Google Play user review (Al-Safoury et al., 2022) for the AtennaPod (Ate, 2023b) open-source app, a podcast manager. The main parts of a user review are the body and rating. The review body contains a textual description of the user feedback about a bug encountered, a new feature requested, etc., expressed without a formal structure (Guzman et al., 2015). For example, in Figure 1(b), a user describes the problem of resuming a podcast after the app disconnects from the car.
A user in this review reported a bug informally. However, this bug has been reported more formally in GitHub issues of the AtennaPod repository (Ate, 2023a). Figure 1(a) shows an issue related to the same bug with the label “type: possible bug”. The main parts of an issue include the title, body, and label. As we can see, a description of the same bug mentioned in the user review (Figure 1(b)) is provided in the Current Behavior section of the body with slightly different constituent words. The issue title also briefly describes this problem. In this issue, other sections provide details for reproducing the bug, which is unrelated to our purpose. Therefore, although GitHub issues more technically report encountered bugs or describe requested features, they can be used for detecting the review’s intention by extracting target information.
3. Proposed Approach
@@
Figure 2. An Approach Overview
3.1. Data Collection
We first collect data needed for subsequent steps of our approach. Therefore, we investigate the Android repositories available on GitHub to identify appropriate ones that (i) contain adequate labelled issues and (ii) provide relevant and accurate data for creating the auxiliary dataset. Liu et al. (Liu et al., 2020) provided the AndroZooOpen dataset, which includes information about more than 46K Android repositories. To our best knowledge, this dataset is the largest available collection containing GitHub repositories related to open-source Android apps. This paper uses the AndroZooOpen dataset to find appropriate repositories by examining extracted information.
In January 2023, approximately 43K repositories of this dataset were available on GitHub. We updated important fields, such as the contributor and star counts, to make our investigation more accurate and get labeled issues using the GitHub REST API (RES, 2023). Most available repositories had no labeled issue, and only 5,258 (12.1%) repositories contained at least one. Figure 3 shows the repository counts for each number of labeled issues in these 5,258 repositories. As can be seen, most of the repositories have a few labeled issues. Although 20% of them contain more than 30 labeled issues, they constitute 92.3% of collected issues. Therefore, repositories with less than 30 labeled issues provide little information. In order to increase the quality and performance of the auxiliary dataset, we eliminate them for two reasons.
First, in preparing the auxiliary dataset, we ignore unrelated labels to our review intentions (Section 3.2.1) and then identify and extract the target information from selected issues (Section 3.2.2). As explained, more than two-thirds of the collected issues, which are irrelevant to our intentions or do not provide sufficient information, are eliminated in these steps. The presence of many repositories with few labeled issues that ultimately provide little helpful information can reduce the performance of finding similar apps (Section 3.4.1). Second, Figure 4 shows the distribution of the number of contributors and stars for repositories with at least one labeled issue. HLI (High-Labeled Issue) and LLI (Low-Labeled Issue) denote repositories with a minimum and a maximum of 30 labeled issues, respectively. As can be seen, more developers are involved in HLI repositories, which causes more monitoring. Also, developers have given HLI repositories more stars. Considering stars as a measure of repository popularity, following Borges et al. (Borges et al., 2016), we can conclude that HLI repositories have garnered more interest from developers. Therefore, data provided by HLI repositories with more monitoring and popularity can be more reliable.
@@ 3.1. Data Collection
Table 1. Preliminary Collected Data
@@
Figure 3. Repository Counts per Each Number of Labelled Issues (Repos=Repositories)
Figure 4. The Comparison of HLI and LLI Repositories
3.2. Preparing the Auxiliary Dataset
@@
Previous studies have classified user reviews according to their intentions (Dąbrowski et al., 2022). They have focused on different perspectives and granularity to assist developers in maintenance tasks (Villarroel et al., 2016; Sorbo et al., 2016). Table 2 displays the target intention classes of reviews in several studies with available labeled datasets. As seen, two intentions, namely Bug Report and Feature Request, have been repeatedly used with slight differences in their naming. Therefore, given the available data required to create the primary dataset, we follow Devine et al. (Devine et al., 2023) and consider the three most frequent classes of Bug Report, Feature Request, and Other as our target review intentions.
Table 2. Target Intention Classes in Previous Studies (*For studies that presented multi-level taxonomy, we only mentioned the first level.)
@@
Pagano et al. (Pagano and Maalej, 2013) community, requirement, rating, user experience *
Maalej et al. (Maalej et al., 2016) bug report, feature request, user experience, rating
Panichella et al. (Panichella et al., 2015) information giving, information seeking, feature request, problem discovery, other
Gu et al. (Gu and Kim, 2015) aspect evaluation, bug reports, feature requests, praise and others
Guzman et al. (Guzman et al., 2015) bug report, feature strength, feature shortcoming, user request, praise, complaint, usage scenario
Ciurumelea et al. (Ciurumelea et al., 2017) compatibility, usage, resources, pricing, and protection*
Jha et al. (Jha and Mahmoud, 2018) bug report, feature request, other
Scalabrino et al. (Scalabrino et al., 2019) functional bug report, suggestion for new feature, report of performance problems, report of security issues, report of excessive energy consumption, request for usability improvements and other
3.2.1. Filtering the Issue Labels
In GitHub, developers assign labels to an issue based on various criteria, such as intentions (e.g., bug, enhancement), states (e.g., duplicate, invalid), development stages (e.g., implementation, design), app components or features (e.g., component: client, feature: history), priorities (e.g., low priority, high priority), etc. These labels are helpful in managing and organizing issues more efficiently. By default, GitHub provides seven labels: Bug, Enhancement, Documentation, Duplicate, Good First Issue, Help Wanted, Invalid, Question, and Wontfix. Developers also have the option to create new labels for categorizing their issues, which we refer to as user-defined labels. There are no specific rules for defining these labels (Izadi et al., 2022). As a result, they can be heterogeneous and diverse. Different labels may represent a similar concept. For instance, labels like Type: Enhancement, Enhancement/Feature, Feature-Request are used in many repositories to represent the same concept as the default label Enhancement. Due to this variability, Izadi et al. (Izadi et al., 2022) manually analyzed the issue labels according to their goals. In the previous phase, we collected 318K labeled issues, comprising 5,641 distinct labels. Figure 5 displays the 20 most frequently occurring labels in our preliminary dataset. Similar labels may vary in form (e.g., bug, Bug) or constituent words (e.g., enhancement, feature request). To address this, we first preprocess the labels to unify their forms and remove noisy characters. Then, we manually identify the semantically relevant labels aligned with our review intentions.
Figure 5. The Most Frequent Labels in Preliminary Dataset
@@ 3.2.1. Filtering the Issue Labels
  4. (4)
    Stemming using the SnowballStemmer function in the NLTK library (NLT, 2023): For example, three labels reproduced, reproduction, and reproducible, which refer to the same concept, are converted to a single term reproduc.
  5. (5)
@@
Table 3. Samples of the Provided Information for Each Preprocessed (Pre) Label
@@
The app crashes, either the foreground or background. **To Reproduce** Steps to reproduce the behavior: (1) Go to Create a new list (2) Click on Ok in the list name dialog, important step (3) Rotate the app
@@
Table 4. Some Preprocessed Labels Related to Target Review Intentions
@@
3.2.2. Extracting Target Information